# RRM: Robust Reward Model Training Mitigates Reward Hacking

### v1

@@ 1 Introduction
Figure 1: The pipeline of our proposed robust reward model (RRM), which aim to decouple contextual preference quality signal and context-free artifacts. Suppose a proportion of chosen responses have certain artifact (bold-face wrapped with “**∗ ∗” in this figure), the reward model can hack the pattern and choose the response with the artifact instead of carefully reading the prompt. With our data augmentations, we can effectively balance the context-free artifacts in chosen and rejected responses, thus ensuring a more robust reward model during inference.
@@ 2 Preliminaries
The second predominant approach is the pairwise ranking model (Zhao et al., 2023; Jiang et al., 2023), which takes a prompt and a pair of responses as the input, and directly predicts the probability $\mathbb{P}(y_{1}\succ y_{2}|x)$, which subsumes the BT model as a subclass. In the literature, the pairwise preference model has shown to outperform pointwise BT reward both empirically (Zhao et al., 2023; Jiang et al., 2023; Dong et al., 2024) and theoretically (Ye et al., 2024) due to its flexibility and larger function class capacity. Specifically, we denote the pairwise ranking model as $\rho_{\psi}(x,y_{1},y_{2})$ and leverage the next token prediction ability of the language model to format the sample as:
@@
Then, the $\rho_{\psi}(x,y_{1},y_{2})$ outputs either “A” or “B” as preferred one. We use the probability of decoding “A” as estimation of the preference probability ^$\hat{{\mathbb{P}}}(y_{1}\succ y_{2}|x)$ [^1] We randomly flip response pairs and associated labels to remove positional bias.. In this work, we use the pairwise ranking model for its superior performance and flexibility.
@@
where $\beta>0$ is the KL penalty coefficient. Several algorithms have been proposed to solve the above optimization, including PPO (Schulman et al., 2017; Ziegler et al., 2019), SLiC (Zhao et al., 2023), DPO (Rafailov et al., 2024), RSO (Liu et al., 2024b), and IPO (Azar et al., 2024). For a stable evaluation process, we use DPO in this work for preference alignment. For a given preference dataset $\mathcal{D}_{\text{p}}$, DPO use the following loss function:
@@
Reward model is not perfect due to its limited model size, limited training data, and distribution shift between training data and alignment prompts and responses (Eisenstein et al., 2023; Gao et al., 2023; Guo et al., 2024; Xiong et al., ). Several works have been proposed to mitigate reward hacking. They mainly focus on observable artifacts such as length (Chen et al., ; Dubois et al., 2024a; Shen et al., 2023b). Shen et al. (2023a) propose to enforce the consistency in reward model via data augmentation. Reward hacking can also be mitigated during policy training with post-adjusted reward (Park et al., 2024) or with post-training model merge (Lin et al., 2023). We focus on improving the reward model by addressing reward hacking from a causal perspective.
@@
Causal inference can be embedded in graphical model frameworks as a directed acyclic graph (DAG) with causal relationship represented as a directed edge (Pearl, 2009; Lee et al., 2020). We say a random vector $X$ to be faithful with respect to a DAG $\mathcal{G}=(\mathcal{V},\mathcal{E})$ if for any $i,j\in\mathcal{V}$, and any subset $S\subseteq V\backslash\{i,j\}$,
@@
where $X^{i}\perp X^{j}\mid X^{S}$ denotes $X^{i}$ and $X^{j}$ are independent conditional on $X^{S}$. The definition of d-separation is as follows: suppose we are given a DAG $\mathcal{G}$; then, for two nodes $i,j\in\mathcal{V}$, a subset $S$ of $V\backslash\{i,j\}$ d-connects $i$ and $j$ if there exists a path $L$ between $i$ and $j$ such that every collider in $L$ either belongs to $S$ or has a descendent in $S$, and no other node in $L$ belongs to $S$. If $S$ does not d-connect $i$ and $j$, then it d-separates $i$ and $j$. See Appendix A.5 for more details.
@@ 3.1 Causal framework
Figure 2: Causal graph of reward model. $X$ is the prompt, $Y_{1},Y_{2}$ are two responses, $S$ is the contextual signal that depends on input prompt and two responses. $A$ is the context-free artifact that only depends on two responses. $C$ is the preference label. Traditional reward model cannot differentiate the two DAGs on whether there is a causal edge from $A$ to $C$. Our work uses the augmented dataset to eliminate the edge from $A$ to $C$.
We formulate a DAG $\mathcal{G}$ to model the causal relationships among different quantities (Figure 2). We assume the distribution of $(X,Y_{1},Y_{2},S,A,C)$ to be faithful to the DAG. $S$ is the sufficient statistic (Lehmann & Casella, 2006) that captures the contextual effect that one response fulfills the need of the prompt better than the other. $A$ is the sufficient statistic that captures the context-free artifacts that only depend on two responses. Such artifacts can include length, format (bold faces, bullet points, markdown, etc), and certain patterns ($n$-grams such as “Sure, here is the response:”). In traditional reward model training, the model may hack the patterns in $(Y_{1},Y_{2})$. Suppose 80% of winning responses to be longer, then the reward model can get 80% accuracy by just counting the number of tokens. Formally, we construct two hypothesis:
@@
The desired behavior of reward model is to ignore the artifact $A$ in determining the preference label $C$, which corresponds to $\mathcal{H}_{0}$. To achieve that, we can utilize two d-separation relationships of the DAG $\mathcal{G}$.
@@
In addition, we can set the preference rule based on the DAG $\mathcal{G}$. We say response $y$ is contextual to $x$ if they are from the same triplet in $\mathcal{D}_{\text{hf}}=\{x^{(i)},y_{w}^{(i)},y_{l}^{(i)}\}_{i=1}^{N}$. For example, $y_{w}^{(i)}$ and $y_{l}^{(i)}$ are contextual to $x^{(i)}$, but $y_{w}^{(j)}$ and $y_{l}^{(j)}$ are not contextual to $x^{(i)}$ for $j\neq i$. Then for $(x,y_{1},y_{2})$, we have the following rules:
@@
3.2 Data augmentation
To fix the issue mentioned in Proposition 3.1, we can effectively utilize R1&R2. Given $\mathcal{D}_{\text{hf}}=\{x^{(i)},y_{w}^{(i)},y_{l}^{(i)}\}_{i=1}^{N}$, we can first expand the dataset as $\tilde{\mathcal{D}}_{\text{hf}}=\{x^{(i)},y_{w}^{(i)},y_{l}^{(i)},x^{(\sigma_{1}(i))},y_{w}^{(\sigma_{1}(i))},y_{l}^{(\sigma_{1}(i))},x^{(\sigma_{2}(i))},y_{w}^{(\sigma_{2}(i))},y_{l}^{(\sigma_{2}(i))}\}_{i=1}^{N}$, where $\sigma_{1}:[N]\rightarrow[N]$ and $\sigma_{2}:[N]\rightarrow[N]$ are two different invertible permutation functions randomly sampled from permutation group $S_{N}$. There are in total $453\times{6\choose 2}=$ 45 possible $(x,y_{1},y_{2})$ unordered triplets from each element in $\tilde{\mathcal{D}}_{\text{hf}}$. This is because there are 3 possible prompts with 2 choices among 6 responses and we treat $(x,y_{1},y_{2})$ and $(x,y_{2},y_{1})$ as the same one.
@@ 3.3 Connection to existing works
This work shows the issues of reward models on the instruction and response consistencies when switching instruction or response to another similar one. It proposes a data augmentation training approach and retrieval-augmented inference technique to improve the consistencies of reward models. On contrary, by considering all possible combinations of $(x,y_{1},y_{2})$ across difference examples, our approach uses the organic data from the dataset, which can effectively eliminate the artifacts existing in the dataset.
@@ 4.1 Settings
We study RRM using the preference dataset curated by RLHFlow [^3] https://huggingface.co/datasets/RLHFlow/pair_preference_model_dataset (Dong et al., 2024), which has been used to train a series of strong open-source preference models as evaluated by the Rewardbench (Lambert et al., 2024). The dataset consists of 700K preference pairs, which is a mixture of HH-RLHF (Bai et al., 2022a), SHP (Ethayarajh et al., 2022), HelpSteer (Wang et al., 2023), PKU-SafeRLHF (Ji et al., 2024), UltraFeedback (Cui et al., 2023), UltraInteract (Yuan et al., 2024), Distilabel-Capybara (Daniele & Suphavadeeprasit, 2023), and Distilabel-Orca (Lian et al., 2023). See Appendix A.2 for details of dataset composition.
@@
We first train a pairwise ranking reward model (RM) from Gemma-2-9b-it. With the augmentation illustrated in Equation 5, we can get 14X additional examples, most of which can be too easy for RM to learn. To reduce the augmented data size, we first conduct inference on random 50% of the augmented data using the trained RM, and leave the examples with |^$|\hat{\mathbb{P}}(A\succ B)-\mathbb{P}^{*}(A\succ B)|\geq 0.2$, where ^$\hat{\mathbb{P}}(A\succ B)$ is winning probability calculated by RM and $\mathbb{P}^{*}(A\succ B)$ is the ground truth probability [^4] The ground truth probability is 1 if A is preferred over B, 0 if B is preferred over A, and 0.5 if they tie.. We get 2.4M training examples by merging the filtered augmented data and original RM training data. Then we use the same training recipe to get the robust reward model (RRM). We train the reward models until convergence using AdamW (Loshchilov, 2017) optimizer with learning rate 1e-6 and batch size 128555We pick the best hyperparameter by grid search..
@@ 4.2 Main Results
We investigate the quality of reward models by evaluating the aligned policies. To study the effect of adding “Neutrals” in Equation 5, we also train a reward model without augmented neutrals (-Neutrals). The results are summarized in Table 2. As expected, ODIN (Chen et al., ) [^6] Training details in Appendix A.4. shows shorter responses than RM and RRM since it explicitly disentangles the length from quality. RRM shows the best performance on both benchmarks over ODIN and RM, with shorter responses generated than RM, suggesting it effectively controls the length as one of the artifact. The added “Neutrals” have slight improvements on first-turn MT-Bench and AlpacaEval-2.
Reward Policy MT-Bench777we do not evaluate BoN policies on MT-Bench because it involves multi-turn. AlpacaEval-2
@@ 4.3 Length Analysis
We study the length (number of tokens) distribution of reward model training datasets. Length is one common artifact that shows bias on both policy training and evaluation. We hypothesize that the bias can possibly come from the reward model training data. The one used in training RM is not well calibrated and the chosen responses are longer on average (Figure 3(a) and 3(b)) and by frequency (Figure 3(c)). On contrary, the RRM training data is better calibrated with length more balanced between chosen and rejected responses in each length bin (Figure 3(a) and 3(b)). We further provide length analysis for each data source in Appendix A.2.
@@ 5 Related Works
The first RLHF framework is based on the proximal policy optimization (PPO) algorithm, which was first popularized by Christiano et al. (2017) and further developed by Bai et al. (2022a); Ouyang et al. (2022). However, getting the deep RL method (PPO) work is challenging especially in the era of LLMs (Choshen et al., ; Engstrom et al., 2020). In recognition of this issue, another line of works proposes direct alignment algorithms, where notable examples include SLiC (Zhao et al., 2023), DPO (Rafailov et al., 2024), IPO (Azar et al., 2024), KTO (Ethayarajh et al., 2024), ORPO (Hong et al., 2024), SimPO (Meng et al., 2024), and DRO (Richemond et al., 2024). These algorithms directly optimize a supervised target to optimize the policy model without constructing a reward model first, hence the name direct alignment algorithms. However, these algorithms learning from a fixed dataset are offline and often off-policy without further exploration of the environment. RSO (Liu et al., 2024b) emphasizes the importance of reward model and fixes the distribution shift problem to improve the DPO training, followed by list-wise alignment (Liu et al., 2024a) and the online (iterative) training frameworks (Xiong et al., ; Guo et al., 2024; Calandriello et al., ). Alternatively, there is also a line of work based on the best-of-n sampling, such as RAFT (Dong et al., ), BOND (Sessa et al., 2024), BoNBoN alignment (Gui et al., 2024). These algorithms leverage a reward model to rank the generated responses and distill knowledge from the best responses. Our approach can benefit RLHF algorithms relying on a reward model.

### v2

@@ 1 Introduction
Figure 1: The pipeline of our proposed robust reward model (RRM), which aims to decouple contextual preference quality signal and context-free artifacts. Suppose a proportion of chosen responses have certain artifact (bold-face wrapped with “**∗ ∗” in this figure), the reward model can hack the pattern and choose the response with the artifact instead of carefully reading the prompt. With our data augmentations, we can effectively balance the context-free artifacts in chosen and rejected responses, thus ensuring a more robust reward model during inference.
@@ 2 Preliminaries
The second predominant approach is the pairwise ranking model (Zhao et al., 2023; Jiang et al., 2023), which takes a prompt and a pair of responses as the input, and directly predicts the probability $\mathbb{P}(y_{1}\succ y_{2}|x)$, which subsumes the BT model as a subclass. In the literature, the pairwise preference model has shown to outperform pointwise BT reward both empirically (Zhao et al., 2023; Jiang et al., 2023; Dong et al., 2024) and theoretically (Ye et al., 2024) due to its flexibility and larger function class capacity. Specifically, we denote the pairwise ranking model and leverage the next token prediction ability of the language model to format the sample as:
@@
Then, the model outputs either “A” or “B” as preferred one. We use the probability of decoding “A” as estimation of the preference probability ^$\hat{{\mathbb{P}}}(y_{1}\succ y_{2}|x)$ [^1] We randomly flip response pairs and associated labels to remove positional bias.. In this work, we use the pairwise ranking model for its superior performance and flexibility.
@@
[…]
@@
Reward model is not perfect due to its limited model size, limited training data, and distribution shift between training data and alignment prompts and responses (Eisenstein et al., 2023; Gao et al., 2023; Guo et al., 2024; Xiong et al., ). Several works have been proposed to mitigate reward hacking. One line of works focus on observable artifacts such as length (Chen et al., ; Dubois et al., 2024a; Shen et al., 2023b). Shen et al. (2023a) propose to enforce the consistency in reward model via data augmentation. To improve generalization, reward model ensembles can mitigate (but do not eliminate) reward hacking (Coste et al., 2023; Eisenstein et al., 2023; Ramé et al., 2024b). Reward hacking can also be mitigated during policy training with post-adjusted reward (Park et al., 2024) or with post-training model merge (Lin et al., 2023). We focus on improving the reward model by addressing reward hacking from a causal perspective.
@@
Causal inference can be embedded in graphical model frameworks as a directed acyclic graph (DAG) $\mathcal{G}=(\mathcal{V},\mathcal{E})$ with variables represented as nodes in $\mathcal{V}$ and causal relationship represented as a directed edge (Pearl, 2009; Lee et al., 2020) in $\mathcal{E}$. We say a random vector $X$ to be faithful with respect to a DAG $\mathcal{G}=(\mathcal{V},\mathcal{E})$ if for any $i,j\in\mathcal{V}$, and any subset $S\subseteq\mathcal{V}\backslash\{i,j\}$,
@@
where $X^{i}\perp X^{j}\mid X^{S}$ denotes $X^{i}$ and $X^{j}$ are independent conditional on $X^{S}$. The “d” in d-separation stands for dependence. Thus if $X^{i}$ and $X^{j}$ are d-separated relative to a set of variables $X^{S}$ in a directed graph, then they are independent conditional on $X^{S}$ in all probability distributions such a graph can represent. The definition of d-separation is as follows: suppose we are given a DAG $\mathcal{G}$; then, for two nodes $i,j\in\mathcal{V}$, a subset $S$ of $\mathcal{V}\backslash\{i,j\}$ d-connects $i$ and $j$ if there exists a path $L$ between $i$ and $j$ such that every collider in $L$ either belongs to $S$ or has a descendent in $S$, and no other node in $L$ belongs to $S$. If $S$ does not d-connect $i$ and $j$, then it d-separates $i$ and $j$. See Appendix A.4 for more details.
@@ 3 Robust Reward Model Training
We first formulate the reward model training in a causal framework, then we augment the reward model training data based on the causal rules.
@@ 3.1 Causal framework
Figure 2: Causal graph of reward model. $X$ is the prompt. $Y_{1},Y_{2}$ are two responses. $S$ is the contextual signal that depends on input prompt and two responses. $A$ is the context-free artifact that only depends on two responses. $C$ is the preference label. Traditional reward model cannot differentiate the two DAGs on whether there is a causal edge from $A$ to $C$. Our work uses the augmented dataset to eliminate the edge from $A$ to $C$.
We formulate a DAG $\mathcal{G}$ to model the causal relationships among different quantities (Figure 2). $X$ is the prompt, and $Y_{1},Y_{2}$ are two responses. $S\in\mathbb{R}$ is the contextual signal that depends on input prompt and two responses. $A\in\mathbb{R}$ is the context-free artifact that only depends on two responses. $C\in\{0,1\}$ is the preference label, where $C=1$ means $Y_{1}$ is preferred over $Y_{2}$ and $C=0$ means the other way around. We assume the distribution of $(X,Y_{1},Y_{2},S,A,C)$ to be faithful to the DAG. We assume the preference label $C$ can be captured by $S$ and $A$, which is to say $\mathbb{P}(C|X,Y_{1},Y_{2})=\mathbb{P}(C|S,A)$. We assume the $S$ to be the sufficient statistic (Lehmann & Casella, 2006) that captures the contextual effect that one response fulfills the need of the prompt better than the other. We assume $A$ to the sufficient statistic that captures the context-free artifacts that only depend on two responses. Such artifacts can include length, format (bold faces, bullet points, markdown, etc), and certain patterns ($n$-grams such as “Sure, here is the response:”). In traditional reward model training, the model may hack the patterns in $(Y_{1},Y_{2})$. Suppose 80% of winning responses to be longer, then the reward model can get 80% accuracy by just counting the number of tokens. Formally, we construct two hypothesis:
@@
Proof.
As an example of the hypotheses being indistinguishable, let’s consider a special case of $A$ and $S$ being perfectly correlated. More formally, assume $S=s(X,Y_{1},Y_{2})+\epsilon_{s}$ with certain non-linear function $s$ and $\epsilon_{s}\sim N(0,\sigma_{s})$, and similarly $A=a(X,Y_{1},Y_{2})+\epsilon_{a}$ with non-linear function $a$ and $\epsilon_{a}\sim N(0,\sigma_{a})$. Suppose $\mathbb{P}(C=1|X,Y_{1},Y_{2})=\sigma(\beta_{s}S+\beta_{a}A+\alpha+\epsilon_{c})$ with constants $\beta_{s},\beta_{a},\alpha\in\mathbb{R}$ and random error $\epsilon_{c}\perp(S,A)$. If $\beta_{a}=0$, then $\mathcal{H}_{0}$ is true. If $\beta_{a}=1$, then $\mathcal{H}_{1}$ is true. In extreme case that $Corr(S,A)=1$, then $A=\beta_{as}S+\alpha_{a}$ for some constants $\alpha_{a}\in\mathbb{R}$ and $\beta_{as}\in\mathbb{R}^{+}$. Then the model cannot tell if $\beta_{a}=0$ or not. This is because when $\beta_{a}=0$, we can still reparametrize it as $\mathbb{P}(C=1|X,Y_{1},Y_{2})=\sigma((\beta_{s}-\beta_{as})S+A+(\alpha-\alpha_{a})+\epsilon_{c})$. ∎
The desired behavior of a reward model is to determine the preference label $C$ ignoring the artifact $A$, which corresponds to $\mathcal{H}_{0}$. To achieve that, we can utilize two d-separation relationships of the DAG $\mathcal{G}$.
@@
3.2 Data augmentation
To fix the issue mentioned in Proposition 3.1, we can effectively utilize R1&R2. In particular, we propose to augment data with by adding the permuted pairs of generated responses.
Possible Combinations
Given the dataset of triplets $\mathcal{D}_{\text{hf}}=\{t^{(i)}\}_{i=1}^{N}$ with $t^{(i)}=(x^{(i)},y_{w}^{(i)},y_{l}^{(i)})$, we can first expand the dataset as $\tilde{\mathcal{D}}_{\text{hf}}=\{t^{(i)},t^{(\sigma_{1}(i))},t^{(\sigma_{2}(i))}\}_{i=1}^{N}$, where $\sigma_{1}:[N]\rightarrow[N]$ and $\sigma_{2}:[N]\rightarrow[N]$ are two different invertible permutation functions randomly sampled from permutation group $S_{N}$. In practice, we can shuffle the dataset twice to achieve $\sigma_{1}$ and $\sigma_{2}$. There are in total $453\times{6\choose 2}=$ 45 possible $(x,y_{1},y_{2})$ unordered triplets from each element in $\tilde{\mathcal{D}}_{\text{hf}}$. This is because there are 3 possible prompts with 2 choices among 6 responses and we treat $(x,y_{1},y_{2})$ and $(x,y_{2},y_{1})$ as the same one.
Preference Labels
For the unordered triplet, we can set the preference rule based on the DAG $\mathcal{G}$. We say response $y$ is contextual to $x$ if they are from the same triplet in $\mathcal{D}_{\text{hf}}=\{x^{(i)},y_{w}^{(i)},y_{l}^{(i)}\}_{i=1}^{N}$. For example, $y_{w}^{(i)}$ and $y_{l}^{(i)}$ are contextual to $x^{(i)}$, but $y_{w}^{(j)}$ and $y_{l}^{(j)}$ are not contextual to $x^{(i)}$ for $j\neq i$. Then for $(x,y_{1},y_{2})$, we have the following rules:
@@ 3.2 Data augmentation
Here we assume that $y_{l}^{(i)}$ is still an acceptable response for $x^{(i)}$ because it is usually generated by a language model conditional on $x^{(i)}$.
Augmented Triplets
@@ 3.3 Connection to existing works
This work shows the issues of reward models on the instruction and response consistencies when switching instruction or response to another similar one. It proposes a data augmentation training approach and retrieval-augmented inference technique to improve the consistencies of reward models. On contrary, by considering all possible combinations of $(x,y_{1},y_{2})$ across different examples, our approach uses the organic data from the dataset, which can effectively eliminate the artifacts existing in the dataset.
@@ 4.1 Settings
We study RRM using the preference dataset curated by RLHFlow [^4] https://huggingface.co/datasets/RLHFlow/pair_preference_model_dataset (Dong et al., 2024), which has been used to train a series of strong open-source preference models as evaluated by the Reward-Bench (Lambert et al., 2024). The dataset consists of 700K preference pairs, which is a mixture of HH-RLHF (Bai et al., 2022a), SHP (Ethayarajh et al., 2022), HelpSteer (Wang et al., 2023), PKU-SafeRLHF (Ji et al., 2024), UltraFeedback (Cui et al., 2023), UltraInteract (Yuan et al., 2024), Distilabel-Capybara (Daniele & Suphavadeeprasit, 2023), and Distilabel-Orca (Lian et al., 2023). We list the data sources and number of examples in Table 1. The authors of the original paper delete the samples with similar scores when the scores are available because when the model is well calibrated, these samples are more likely to mislabelled. Thus the total number is smaller than the sum of each individual datasets.
Source Number of Examples
Table 1: Composition of reward model training dataset
@@
We first train a pairwise ranking reward model (RM) from Gemma-2-9b-it. With the augmentation illustrated in Equation 5, we can get 14X additional examples, most of which can be too easy for RM to learn. To reduce the augmented data size, we first conduct inference on random 50% of the augmented data using the trained RM, and leave the examples with |^$|\hat{\mathbb{P}}(A\succ B)-\mathbb{P}^{*}(A\succ B)|\geq 0.2$, where ^$\hat{\mathbb{P}}(A\succ B)$ is winning probability calculated by RM and $\mathbb{P}^{*}(A\succ B)$ is the ground truth probability [^13] The ground truth probability is 1 if A is preferred over B, 0 if B is preferred over A, and 0.5 if they tie.. We get 2.4M training examples by merging the filtered augmented data and original RM training data. Then we use the same training recipe to get the robust reward model (RRM). We train the reward models for 1 epoch using AdamW (Loshchilov, 2017) optimizer with learning rate 1e-6 and batch size 128141414We found 1 epoch is best for reward model training and we pick the best hyperparameter by grid search..
@@ 4.2 Main Results
We investigate the quality of reward models by evaluating the aligned policies. To study the effect of adding “Neutrals” in Equation 5, we also train a reward model without augmented neutrals (-Neutrals). The results are summarized in Table 3. As expected, ODIN (Chen et al., ) [^15] Training details in Appendix A.3. shows shorter responses than RM and RRM since it explicitly disentangles the length from quality. RRM shows the best performance on MT-Bench first turn and AlpacaEval-2 over ODIN and RM, with shorter responses generated than RM, suggesting it effectively controls the length as one of the artifact. The added “Neutrals” have slight improvements on first-turn MT-Bench and AlpacaEval-2.
Reward Policy MT-Bench161616we do not evaluate BoN policies on MT-Bench because it involves multi-turn. AlpacaEval-2
@@ 4.3 Length Analysis
We study the length (number of tokens) distribution of reward model training datasets. Length is one common artifact that shows bias on both policy training and evaluation. We hypothesize that the bias can possibly come from the reward model training data. The one used in training RM is not well calibrated and the chosen responses are longer on average (Figure 3(a)) and by frequency (Figure 3(c)). On contrary, the RRM training data is better calibrated with length more balanced between chosen and rejected responses in each length bin (Figure 3(b) and 3(c)). We further provide length analysis for each data source in Appendix A.1.
@@ 5 Related Works
The first RLHF framework (Stiennon et al., 2020) is based on the proximal policy optimization (PPO) algorithm, which was first popularized by Christiano et al. (2017) and further developed by Bai et al. (2022a); Ouyang et al. (2022). However, getting PPO work is challenging especially in the era of LLMs (Choshen et al., ; Engstrom et al., 2020). In recognition of this issue, another line of works propose direct alignment algorithms, where notable examples include SLiC (Zhao et al., 2023), DPO (Rafailov et al., 2024), IPO (Azar et al., 2024), KTO (Ethayarajh et al., 2024), ORPO (Hong et al., 2024), SimPO (Meng et al., 2024), and DRO (Richemond et al., 2024). These algorithms directly optimize a supervised target to optimize the policy model without constructing a reward model first, hence the name direct alignment algorithms. However, these algorithms learning from a fixed dataset are offline and often off-policy without further exploration of the environment. RSO (Liu et al., 2024b) emphasizes the importance of reward model and fixes the distribution shift problem to improve the DPO training, followed by list-wise alignment (Liu et al., 2024a) and the online (iterative) training frameworks (Xiong et al., ; Guo et al., 2024; Calandriello et al., ). Alternatively, there is also a line of work based on the best-of-n sampling, such as RAFT (Dong et al., ), BOND (Sessa et al., 2024), BoNBoN alignment (Gui et al., 2024). These algorithms leverage a reward model to rank the generated responses and distill knowledge from the best responses. Our approach can benefit RLHF algorithms relying on a reward model.